# Modeling Maritime Transportation Behavior Using AIS Trajectories and Markovian Processes in the Gulf of St. Lawrence

### v2

@@ III-A Trajectory Representation
[…]
@@ III-B State-Space Definition
We transform continuous trajectories into discrete sequences by partitioning the maritime domain into a finite set of states using a uniform hexagonal grid. In comparison to conventional square tiling, hexagonal binning provides uniform neighbor connectivity and reduces directional bias [30], improving spatial fidelity in mobility modeling. To define the spatial state space, we use the hierarchical hexagonal indexing system H3222https://h3geo.org/, which divides the Earth’s surface into equally spaced hexagonal cells [31]. For mobility and dwell-time computations, we select resolution level 6, which produces approximately $8{,}687$ hexagonal cells, each covering around $36\,\mathrm{km}^{2}$.
@@ III-E1 Dataset Segmentation
[…]
@@ III-E3 Metrics Quantization
Many stochastic metrics follow heavy-tailed distributions. To reduce the influence of outliers, we first clip raw node-weight distributions $\mathcal{D}_{w}$ to the $[1^{\text{st}},98^{\text{th}}]$ percentile range. We then apply a spline transformation $\mathfrak{S}$ to enhance visual contrast in low and mid-value ranges across vessel types. In this case, we consider $Q_{\text{low}}$ and $Q_{\text{high}}$ as the percentile thresholds per vessel type over all windows.

### v3

@@ III-A Trajectory Representation
where $(x_{t},y_{t})$ is the latitude and longitude recorded at timestamp $t$, and $T=\{t_{1}<t_{2}<\dots<t_{|T|}\}$ is the set of observation timestamps. Due to irregular AIS transmission intervals, raw trajectories exhibit variable temporal resolution [29]. To standardize trajectories for comparative analysis, we segment each trajectory into 3-hour windows and resample positional data at fixed time intervals $\Delta t$ (set to $1~\text{min}$) using linear interpolation. The standardized trajectory is then given by:
@@ III-B State-Space Definition
We transform continuous trajectories into discrete sequences by partitioning the maritime domain into a finite set of states using a uniform hexagonal grid. In comparison to conventional square tiling, hexagonal binning provides uniform neighbor connectivity and reduces directional bias [30], improving spatial fidelity in mobility modeling. To define the spatial state space, we use the hierarchical hexagonal indexing system H3222https://h3geo.org/, which divides the Earth’s surface into equally spaced hexagonal cells [31]. For mobility and dwell-time computations, we select resolution level 6, which produces approximately 8,6878{,}687 hexagonal cells, each covering around $36\,\mathrm{km}^{2}$.
@@ III-E1 Dataset Segmentation
$\mathcal{T}_{\text{pre}}$ $=\bigl{[}\text{2019‑01‑01},\,\text{2019‑12‑31}\bigr{]},$ (14)
$\mathcal{T}_{\text{pandemic}_{P1}}$ $=\bigl{[}\text{2020‑01‑01},\,\text{2020‑12‑31}\bigr{]},$
$\mathcal{T}_{\text{pandemic}_{P2}}$ $=\bigl{[}\text{2021‑01‑01},\,\text{2021‑12‑31}\bigr{]},$
$\mathcal{T}_{\text{post}}$ $=\bigl{[}\text{2022‑01‑01},\,\text{2022‑12‑31}\bigr{]}.$
@@ III-E3 Metrics Quantization
Many stochastic metrics follow heavy-tailed distributions. To reduce the influence of outliers, we first clip raw node-weight distributions $\mathcal{D}_{w}$ to the $1^{\text{st}},98^{\text{th}}]$ percentile range. We then apply a spline transformation $\mathfrak{S}$ to enhance visual contrast in low and mid-value ranges across vessel types. In this case, we consider $Q_{\text{low}}$ and $Q_{\text{high}}$ as the percentile thresholds per vessel type over all windows.